Which pair of ratios does not form a proportion?
A. 4/10, 18/45
B. 26/2,52/4
C.-2/5, -8/25
D. -4/7, -20/35
Answ. -4/7, -20/35 Right????
: Which pair of ratios does not form a proportion?
A. 4/10, 18/45
B. 26/2,52/4
C.-2/5, -8/25
D. -4/7, -20/35
Answ. -4/7, -20/35 Right????

Answer by checkley77(3639)

answer d) is incorrect. (-4/7)*5=-20/35
answer c) is not an appropriate ratio: (-2/5)*4=-8/20